OpenStudy (anonymous):

Find the general indefinite integral. (4x^(3)-5squareroot x)/x How do you find antiderivatives that involve a quotient?

hartnn (hartnn):

in this case, you just distribute (divide) the 'x' in the denominator to all terms of numerator.

hartnn (hartnn):

what is 4x^3/x =... ? -5 \(\sqrt x\)/x =.... ?

OpenStudy (anonymous):

would it be 4x^2 and -5/x^1/2?

hartnn (hartnn):

yes, both are correct :) now you can integrate them individually , can you ?

OpenStudy (anonymous):

4/3 x^3 but not sure about the other one

hartnn (hartnn):

you apply same formula, 1/ x^1/2 = x^ (-1/2) so here n = -1/2

OpenStudy (anonymous):

so would it be 10x^1/2?

hartnn (hartnn):

+10x^1/2 or -10x^1/2 ??

OpenStudy (anonymous):

+

OpenStudy (anonymous):

why would it be -?

hartnn (hartnn):

let me check..... -5 x^(-1/2) ---> -5 x(-1/2+1) / [-1/2+1] ---> -5 x^(+1/2) / [+1/2] --->-10x^(1/2)

hartnn (hartnn):

the formula has x^(n+1)/ (n+1) right ? so it'll be -

OpenStudy (anonymous):

ohhhhhh yeah it has to be

OpenStudy (anonymous):

so would the answer be 4/3 x^3-10x^(1/2) if you put it together?

hartnn (hartnn):

yes, and add a +c to the end

OpenStudy (anonymous):

got it, thanks for helping me work through that because I had no idea how to start!

hartnn (hartnn):

you are most welcome ^_^